OpenStudy (anonymous):

Choose the equation of the line passing through the point (-3, 1) and parallel to y = -x - 4.

OpenStudy (anonymous):

i got y=x+2

OpenStudy (anonymous):

@sheelo_mughal

OpenStudy (matt101):

Your equation is going to be in y = mx + b form. The point gives us x and y, and the equation gives us m (parallel lines have the same slope, in this case -1). Let's plug in those values and solve for b! y = mx + b 1 = (-1)(-3) + b 1 = 3 + b b = -2 This means your general equation is y = -x-2

OpenStudy (anonymous):

can you check 3 more?

OpenStudy (matt101):

Sure

OpenStudy (anonymous):

The tables below show the values of y corresponding to different values of x: Table A x 3 2 1 y 1 0 0 Table B x 3 5 4 y -2 1 1 Which statement is true for the tables? Both Table A and Table B represent functions. Both Table A and Table B do not represent functions. Table A does not represent a function, but Table B represents a function. Table A represents a function, but Table B does not represent a function.

OpenStudy (anonymous):

I got D

OpenStudy (matt101):

A function refers to a graph where every x-value has exactly 1 y-value. This means different x-values could have the same y-value, as long as there is just one (like what you get in a parabola). This means A is the correct answer.